OpenStudy (cenaida):

Solve the following system of equations by elimination: 3x + 2y= 8 x + y= 2. A. x= -4 y= -2 B. x= 4 y= -2 C. x= 4 y= 1 D. x= 6 y= 3

OpenStudy (anonymous):

Are those two different equations? As in 3x+2y=8 and x+y=2

OpenStudy (cenaida):

yea i think so

OpenStudy (solomonzelman):

3x + 2y= 8 x + y= 2 ---> times (-3) ---> -3x-3y=-6 So, 3x + 2y= 8 -3x-3y=-6 --------- ADD 0x-y=2 y=-2

OpenStudy (solomonzelman):

Solve for x by plugging in y. x + y= 2 x + (-2)= 2 x=4

OpenStudy (solomonzelman):

makes sense?

OpenStudy (cenaida):

yesss then its B thanks again!!

OpenStudy (solomonzelman):

Anytime, I wish all subjects were same hard as math....

OpenStudy (wolf1728):

A) 3x + 2y= 8 B) x + y= 2 Multiply Equation B by -2 B) -2x -2y = -4 Adding this to A) 3x + 2y = 8 x = 4 putting this in equation B 4 + y = 2 y = -2

OpenStudy (solomonzelman):

Yes same thing, you are eliminating y I eliminated x.

OpenStudy (wolf1728):

I always like it when I am working on a problem, along with someone else, that the answers come out exactly the same. All right!!